# From User Surveys to Telemetry-Driven Agents: Exploring the Potential of Personalized Productivity Solutions

### v1

@@ Abstract.
We present a comprehensive, user-centric approach to understand preferences in AI-based productivity agents and develop personalized solutions tailored to users’ needs. Utilizing a two-phase method, we first conducted a survey with 363 participants, exploring various aspects of productivity, communication style, agent approach, personality traits, personalization, and privacy. Drawing on the survey insights, we developed a GPT-4 powered personalized productivity agent that utilizes telemetry data gathered via Viva Insights from information workers to provide tailored assistance. We compared its performance with alternative productivity-assistive tools, such as dashboard and narrative, in a study involving 40 participants. Our findings highlight the importance of user-centric design, adaptability, and the balance between personalization and privacy in AI-assisted productivity tools. By building on the insights distilled from our study, we believe that our work can enable and guide future research to further enhance productivity solutions, ultimately leading to optimized efficiency and user experiences for information workers.
@@
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in ubiquitous and mobile computing
@@ 1. Introduction
In today’s dynamic work environment, productivity plays a pivotal role in determining professional success. Organizations and individuals continuously strive to enhance their efficiency, improve time management, and optimize their work processes. The rapid advancements in artificial intelligence (AI), natural language processing, and machine learning technologies have introduced new opportunities to develop tools that can assist in boosting productivity. Several studies have explored the potential of AI-based tools in various professional contexts, such as project management (Taboada et al., 2023; Dam et al., 2019), team collaboration (Bansal et al., 2020), and decision-making (Jarrahi, 2018). These studies have demonstrated the potential of AI to not only automate routine tasks but also to provide intelligent insights and recommendations that can help users optimize their work processes and achieve their goals more efficiently. To develop a truly effective productivity agent, it is crucial to understand user preferences and the factors that influence productivity. This paper presents a comprehensive two-phase study aimed at investigating user preferences for productivity agents and devising personalized solutions that cater to these preferences. In the first phase, we conducted a user survey to identify the ideal persona for a productivity agent by examining people’s expectations and preferences concerning its features and behaviors, exploring various aspects such as productivity, communication style, agent approach, personality traits, personalization, and privacy. Our findings revealed users’ specific needs and areas where a personalized productivity agent could offer the most significant benefits. We also found that striking a balance between personalization and privacy is essential, as well as the need for adaptability in tone and approach depending on the situation and user preferences.
@@
In the second phase, we developed an AI-based productivity agent that leverages insights from the initial user survey and uses the state-of-the-art GPT-4 model (OpenAI, 2023a). GPT-4 is a powerful Large Language Model (LLM) (Naveed et al., 2023) capable of generating human-like text and understanding context, making it a reasonable choice for creating a personalized productivity agent. To facilitate personalization, information workers participating in our study provided their workplace telemetry data, such as email and meeting habits, to the productivity agent through a custom application that we developed. The agent used the shared high-level telemetry data to tailor its recommendations and insights based on each user’s unique work habits and patterns, providing personalized assistance for a more effective and customized user experience. In a within-subjects study, we then evaluated the AI agent’s performance and compared it with alternative methods of telemetry data presentation, such as dashboard visualization and personalized narratives. This analysis allowed us to better understand user preferences and identify areas for improvement and refinement.
@@
Our objective in this comprehensive study was to delve into the nuanced requirements and preferences of users regarding AI-driven productivity tools. To structure our investigation, we formulated specific research questions for each phase of the study. These questions guided our research methodology and helped us focus on key areas of user interest and technology application:
@@
  • (RQ3): How effective is the AI productivity agent in understanding and addressing users’ productivity challenges based on their telemetry data? This question delves into the AI productivity agent’s performance, particularly focusing on its capability to analyze and respond to individual user productivity needs using the gathered telemetry data.
@@
In addressing the research questions discussed above, we make the following contributions:
  • We conducted a comprehensive user survey with N=363 information workers that revealed users’ specific needs and preferences for productivity agents. For example, we found that the majority of the participants preferred the agent to provide proactive suggestions to enhance productivity and that they were comfortable sharing high-level telemetry data for enhancing the response provided by the productivity agent.
  • We developed a custom application that facilitated user telemetry data sharing, enabling the AI-based productivity agent to offer personalized assistance. The application would scrape telemetry data off of web pages. We utilize the GPT-4 model for extracting information from the web page as well to create the personalized AI productivity agent.
  • We conducted a thorough evaluation of the AI agent’s performance with N=40 participants, highlighting its ability to provide personalized insights and recommendations based on user telemetry data. In addition, we compared the AI agent with alternative methods of telemetry data presentation, indicating user preferences for a combination of dashboard visualization and interactive chat agents.
As we continue to witness rapid advancements in AI, natural language processing, and machine learning technologies, there is a growing potential for further enhancing and refining productivity agents. The incorporation of LLMs in the development of our AI-based productivity agent demonstrates the power of language models in creating personalized, context-aware, and potentially efficient productivity-enhancing tools. Our findings underscore the importance of a user-centric approach to the design and development of AI-based productivity agents and highlight the potential of incorporating design principles, such as semantic zooming (Shneiderman, [n. d.]) and progressive disclosure (Nielsen, 2023), to enhance their effectiveness and user experience. By building on the insights gleaned from this study, future research can continue to refine and optimize productivity-enhancing tools and solutions, ultimately leading to improved efficiency and user-centric experiences for information workers. This paper makes meaningful contributions to the field of HCI and AI-assisted productivity, and we hope these can meaningfully guide practitioners in the development of effective, user-centric solutions in this domain.
@@ 2. Related Work
The landscape of productivity tools has undergone a significant evolution in recent years. The field has progressed from rudimentary task managers and calendar applications to more sophisticated AI-driven virtual assistants, such as Microsoft’s Cortana, Apple’s Siri, and Google’s Assistant. While numerous studies have emphasized the effectiveness of these general-purpose virtual assistants in enhancing user efficiency and satisfaction across various contexts and domains (de Barcelos Silva et al., 2020; Kiseleva et al., 2016; Brill et al., 2019), there remains a gap in exploring AI-driven productivity agents tailored specifically to individual needs. This gap becomes even more pronounced when considering the unique demands of information workers, who face challenges that often diverge from those of the broader workforce or consumers.
@@
Information workers, who primarily engage in data-centric tasks, often require more specialized tools to address their unique professional demands. While general-purpose virtual assistants offer some degree of support, they may not fully cater to the distinct requirements of these professionals such as scheduling meetings, summarizing research reviews, and preparing presentations or reports. In recent years, customized solutions—particularly conversational agents—have emerged as a promising alternative due to their adaptability to individual professional needs (Diederich et al., 2022; Feng et al., 2020). However, the majority of existing research on conversational or chat agents is centered on customer service (Feine et al., 2019; Cui et al., 2017) or marketing and sales (Vaccaro et al., 2018; Van den Broeck et al., 2019). Other common areas of study for agents include education (Hayashi, 2013) and health (Bhirud et al., 2019; Bell et al., 2019). Despite this focus, some progress has been made in researching solutions for information workers. For example, various conversational agent-based approaches have been proposed to assist with scheduling and managing tasks (Toxtli et al., 2018; Gil and Ratnakar, 2008), managing distractions (Tseng et al., 2019; Mark et al., 2018), coordinating meetings (Cranshaw et al., 2017), work reflections (Kocielnik et al., 2018) and even augmenting group decision-making (Shamekhi et al., 2018). Notably, the work by  Grover et al. (2020) demonstrated the design of a chatbot-like agent, complemented by an emotionally expressive video avatar, that aimed to enhance productivity. Their productivity agent assisted N=40 users in scheduling focused tasks, monitoring distractions, and reflecting on their daily mood and goals through a standalone application. Authors reported that the agents led to an increase in scheduled time for focused tasks and that the users felt more satisfied and productive with the agents. Similarly,  Kimani et al. (2019) introduced Amber, a conversational agent designed to support information workers across a diverse range of tasks, such as scheduling, prioritizing tasks, task switching, providing break reminders, managing social media distractions, and reflecting on daily accomplishments. The researchers conducted a field study with 24 participants over six days and found promising results for the potential use of conversational agents in enhancing workplace productivity and well-being. A critical aspect of designing these agents is understanding user preferences. For example, Kimani et al. (2019) conducted an online survey with N=70 participants to investigate when information workers might engage with a conversational agent. The authors found that users needed help with reminders, scheduling tasks, and managing distractions. Another study by Ahire et al. (2022) explored the expectations of N=14 knowledge workers from such agents, with findings highlighting the importance of features related to scheduling, distraction management, and well-being. Khaokaew et al. (2022) conducted a user study of N=40 workers over four weeks to identify user needs for not just conversational agents but for all kinds of digital assistants (DA). The authors report that the participants imagined a DA that supported them by managing their (1) time, (2) tasks, and (3) information. While these insights provide valuable information, there is still a pressing need for a more comprehensive exploration of the specific requirements of information workers. Previous studies have broadly addressed productivity tools without honing in on the unique challenges and needs of information workers. Our research aims to fill this gap by focusing specifically on how these professionals interact with and benefit from AI-driven productivity assistants. Note, in this study, ‘information workers’ refer to professionals whose primary role involves acquiring, manipulating, and generating information (Kuhlthau, 1999).
@@
Recent advancements in natural language processing, particularly the emergence of advanced language models such as GPT-3 and GPT-4, have significantly expanded the potential of AI-driven conversational agents. However, creating agents tailored to individual users’ work habits and patterns remains relatively uncharted territory. Although earlier applications of conversational agents in AI productivity tools, as developed by  Grover et al. (2020) and  Kimani et al. (2019), utilized telemetry or sensing data for customization, it has yet to be thoroughly explored, particularly with the latest large language model-based chat agents that offer a more flexible and powerful alternative to their predecessors due to their semantic understanding. Telemetry data, which provides valuable insights into user behaviors and preferences, is essential for achieving meaningful personalization. Despite its potential to enhance AI-driven productivity agents, the implementation of telemetry data in this context has been limited. Our research aims to bridge these gaps by conducting a comprehensive two-phase study that examines user preferences and incorporates telemetry data for personalization in AI-driven productivity agents. Our approach goes beyond the existing literature by offering an in-depth analysis of both user preferences and the affordances of telemetry data. We delve deeper into its potential by integrating it with advanced AI capabilities, offering a more nuanced and personalized user experience. The use of large language models (LLMs) brings enhanced capabilities in understanding and processing complex user data, providing more accurate and contextually relevant productivity insights. Thus, our study’s innovation lies in the use of cutting-edge LLMs and how we leverage these models to interpret telemetry data, creating a personalized and dynamic productivity tool tailored to individual user needs. We have the position that our findings can pave the way for future research and contribute to the development of more sophisticated productivity tools tailored to the distinct challenges and needs of information workers.

### v2

# From User Surveys to Telemetry-Driven AI Agents: Exploring the Potential of Personalized Productivity Solutions

@@ Abstract.
Information workers increasingly struggle with productivity challenges in modern workplaces, facing difficulties in managing time and effectively utilizing workplace analytics data for behavioral improvement. Despite the availability of productivity metrics through enterprise tools, workers often fail to translate this data into actionable insights. We present a comprehensive, user-centric approach to address these challenges through AI-based productivity agents tailored to users’ needs. Utilizing a two-phase method, we first conducted a survey with 363 participants, exploring various aspects of productivity, communication style, agent approach, personality traits, personalization, and privacy. Drawing on the survey insights, we developed a GPT-4 powered personalized productivity agent that utilizes telemetry data gathered via Viva Insights from information workers to provide tailored assistance. We compared its performance with alternative productivity-assistive tools, such as dashboard and narrative, in a study involving 40 participants. Our findings highlight the importance of user-centric design, adaptability, and the balance between personalization and privacy in AI-assisted productivity tools. By building on these insights, our work provides important guidance for developing more effective productivity solutions, ultimately leading to optimized efficiency and user experiences for information workers.
@@
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in ubiquitous and mobile computing
This paper has been accepted for publication at CSCW 2025 after undergoing peer-review; publication is currently underway.
@@ 1. Introduction
Productivity plays a pivotal role in determining professional success. Organizations and individuals continuously strive to enhance their efficiency, improve time management, and optimize their work processes. The importance of effective productivity tools has become increasingly critical, particularly as studies reveal concerning trends: information workers spend an average of 28% of their workweek on email, 45% of professionals report that excessive meeting time reduces their overall productivity, and 85% of leaders express difficulty in confidently assessing employee productivity in hybrid work settings (Chui et al., 2012; Calendly, [n. d.]; Microsoft, 2022). Despite having access to productivity data through workplace analytics tools, many workers struggle to translate this information into actionable improvements, highlighting the need for more effective, personalized productivity solutions. The rapid advancements in Artificial Intelligence (AI), natural language processing, and machine learning technologies have introduced promising opportunities to address these challenges. Several studies have explored the potential of AI-based tools in various professional contexts, such as project management (Taboada et al., 2023; Dam et al., 2019), team collaboration (Bansal et al., 2020), and decision-making (Jarrahi, 2018). These studies have demonstrated AI’s potential to not only automate routine tasks but also to provide intelligent insights and recommendations that can help users optimize their work processes and achieve their goals more efficiently.
@@
To develop a truly effective productivity agent, it is crucial to understand user preferences and the factors that influence productivity. This paper presents a comprehensive two-phase study aimed at investigating user preferences for productivity agents and devising personalized solutions that cater to these preferences. In the first phase, we conducted a user survey to identify the ideal persona for a productivity agent by examining people’s expectations and preferences concerning its features and behaviors, exploring various aspects such as productivity, communication style, agent approach, personality traits, personalization, and privacy. Our findings revealed users’ specific needs and areas where a personalized productivity agent could offer the most significant benefits. We also found that striking a balance between personalization and privacy is essential, as well as the need for adaptability in tone and approach depending on the situation and user preferences.
@@
In the second phase, we developed an AI-based productivity agent that leverages insights from the initial user survey, telemetry data from users and uses the state-of-the-art GPT-4 model (OpenAI, 2023a). GPT-4 is a powerful Large Language Model (LLM) (Naveed et al., 2023) capable of generating human-like text and understanding context. At the time of this study, GPT-4 represented the state-of-the-art in language models, making it well-suited for providing personalized and contextually relevant productivity advice based on individual user data. We utilized the enterprise version of GPT-4, which functions purely as an inference engine, processing only high-level telemetry data without storing or using any information for model training. Note that our approach to productivity tools is grounded in personal informatics– using productivity tracking systems to support self-reflection and behavior change (Kross et al., 2023). Personal informatics systems, which help individuals collect and reflect on personal data about their behaviors, activities, and habits, have been shown to support behavior change and goal achievement across various domains (Nepal et al., 2022; Li et al., 2010, 2011; Choe et al., 2014). In the context of workplace productivity, personal informatics tools can help workers understand their work patterns, identify areas for improvement, and make informed decisions about changing their behaviors (Meyer et al., 2017; Ekhtiar et al., 2023; Epstein et al., 2016; Das Swain et al., 2023b; Suh et al., 2023). The effectiveness of these tools often depends on how they present information to users and support different stages of behavior change, from initial awareness to sustained action. We explored three distinct approaches to presenting productivity data. Each approach aligns with stages of the personal informatics model: a dashboard visualization supporting reflection through visual pattern exploration, a personalized narrative aiding integration and reflection by connecting data points into a coherent story, and an interactive chatbot agent facilitating action and reflection through data-driven guidance. To personalize the productivity agent, information workers participating in our study provided their workplace telemetry data, such as email and meeting habits, to the agent through a custom application that we developed. The agent used the shared high-level telemetry data to tailor its recommendations and insights based on each user’s unique work habits and patterns, providing personalized assistance for a more effective and customized user experience. In a within-subjects study, we then evaluated the AI agent’s performance and compared it with dashboard visualization and personalized narratives. This analysis allowed us to better understand user preferences and identify areas for improvement and refinement.
@@
Our objective in this comprehensive study was to understand the nuanced requirements and preferences of users regarding AI-driven productivity tools. As such, we focus on user preferences and subjective experiences rather than measuring productivity improvements quantitatively. To structure our investigation, we formulated specific research questions for each phase of the study. These questions guided our research methodology and helped us focus on key areas of user interest and technology application:
@@
  • (RQ3): How effective is the AI productivity agent in understanding and addressing users’ productivity challenges based on their telemetry data? This question centers on the AI productivity agent’s performance, particularly focusing on its capability to analyze and respond to individual user productivity needs using the gathered telemetry data.
@@
In addressing the research questions discussed above, we make several novel contributions:
  • We provide the first systematic comparison of different personal informatics approaches (dashboard, narrative, and interactive agent) for workplace productivity enhancement using AI and workplace telemetry data.
  • Unlike previous studies that focused on general virtual assistants or limited aspects of productivity tools, our large-scale survey (N=363) provides comprehensive insights into user preferences specifically for AI-based productivity agents. For example, we found that the majority of the participants preferred the agent to provide proactive suggestions to enhance productivity and that they were comfortable sharing high-level telemetry data for enhancing the response provided by the productivity agent.
  • We demonstrate a novel application of LLMs in interpreting workplace telemetry data for personalized productivity guidance, showing how advanced language models can transform generic productivity metrics into personalized, actionable insights. We conducted a thorough evaluation of the AI agent’s performance with N=40 participants, highlighting its ability to provide personalized insights and recommendations based on user telemetry data.
  • Our findings reveal important trade-offs between different approaches to presenting productivity data - for instance, while participants strongly preferred dashboards for initial data exploration (55% ranking it first), they found the interactive agent more effective for obtaining personalized recommendations and engaging with their productivity data in depth (showing 3-5x longer engagement times). These empirically-derived insights contribute new knowledge to guide the design of future productivity tools.
As we continue to witness rapid advancements in AI, natural language processing, and machine learning technologies, there is a growing potential for further enhancing and refining productivity agents. The incorporation of LLMs in the development of our AI-based productivity agent demonstrates the power of language models in creating personalized, context-aware, and potentially efficient productivity-enhancing tools. Our findings underscore the importance of a user-centric approach to the design and development of AI-based productivity agents and highlight the potential of incorporating design principles, such as semantic zooming (Shneiderman, 1996) and progressive disclosure (Nielsen, 2023), to enhance their effectiveness and user experience. By building on the insights obtained from this study, future research can continue to refine and optimize productivity-enhancing tools and solutions, ultimately leading to improved efficiency and user-centric experiences for information workers. This paper makes meaningful contributions to the field of HCI and AI-assisted productivity, and we hope these can meaningfully guide practitioners in the development of effective, user-centric solutions in this domain.
@@ 2. Related Work
The landscape of productivity tools has undergone a significant evolution in recent years. The field has progressed from rudimentary task managers and calendar applications to more sophisticated AI-driven virtual assistants, such as Microsoft Cortana (now rebranded as Microsoft Copilot), Siri, and Google Assistant. Although numerous studies have emphasized the effectiveness of these general-purpose virtual assistants in improving user efficiency and satisfaction in various contexts and domains (de Barcelos Silva et al., 2020; Kiseleva et al., 2016; Brill et al., 2019), there remains a gap in exploring AI-driven productivity agents tailored specifically to individual needs. This gap becomes even more pronounced when considering the unique demands of information workers, who face challenges that often diverge from those of the broader workforce or consumers.
@@
Information workers, who primarily engage in data-centric tasks, often require more specialized tools to address their unique professional demands. Although general-purpose virtual assistants offer some degree of support, they may not fully meet the distinct requirements of these professionals, such as scheduling meetings, summarizing research reviews, and preparing presentations or reports. In recent years, customized solutions—particularly conversational agents—have emerged as a promising alternative due to their adaptability to individual professional needs (Diederich et al., 2022; Feng et al., 2020). However, most of the existing research on chat or conversational agents is focused on customer service (Feine et al., 2019; Cui et al., 2017) or marketing and sales (Vaccaro et al., 2018; Van den Broeck et al., 2019). Other common areas of study for agents include education (Hayashi, 2013) and health (Bhirud et al., 2019; Bell et al., 2019). Despite this focus, some progress has been made in researching solutions for information workers. For example, various conversational agent-based approaches have been proposed to assist with scheduling and managing tasks (Toxtli et al., 2018; Gil and Ratnakar, 2008), managing distractions (Tseng et al., 2019; Mark et al., 2018), coordinating meetings (Cranshaw et al., 2017), work reflections (Kocielnik et al., 2018) and even augmenting group decision-making (Shamekhi et al., 2018). Notably, the work by  Grover et al. (2020) demonstrated the design of a chatbot-like agent, complemented by an emotionally expressive video avatar, that aimed to enhance productivity. Their productivity agent assisted N=40 users in scheduling focused tasks, monitoring distractions, and reflecting on their daily mood and goals through a standalone application. The authors reported that the agent led to an increase in scheduled time for focused tasks, and that the users felt more satisfied and productive with the agents. Similarly,  Kimani et al. (2019) introduced Amber, a conversational agent designed to support information workers across a diverse range of tasks, such as scheduling, prioritizing tasks, task switching, providing break reminders, managing social media distractions, and reflecting on daily accomplishments. The researchers conducted a field study with 24 participants over six days and found promising results for the potential use of conversational agents in enhancing workplace productivity and well-being. A critical aspect of designing these agents is understanding user preferences. For example, Kimani et al. (2019) conducted an online survey with N=70 participants to investigate when information workers might engage with a conversational agent. The authors found that users needed help with reminders, scheduling tasks, and managing distractions. Another study by Ahire et al. (2022) explored the expectations of N=14 knowledge workers from such agents, with findings highlighting the importance of features related to scheduling, distraction management, and well-being. Khaokaew et al. (2022) conducted a user study of N=40 workers over four weeks to identify user needs for not just conversational agents but for all kinds of digital assistants (DA). The authors reported that the participants imagined a DA that supported them by managing their (1) time, (2) tasks, and (3) information. While these insights provide valuable information, there is still a pressing need for a more comprehensive exploration of the specific requirements of information workers. Previous studies have broadly addressed productivity tools without honing in on the unique challenges and needs of information workers. Our research aims to fill this gap by focusing specifically on how these professionals interact with and benefit from AI-driven productivity assistants. Note, in this study, ‘information workers’ refer to professionals whose primary role involves acquiring, manipulating, and generating information (Kuhlthau, 1999).
The emergence of advanced language models such as GPT-3 and GPT-4, have significantly expanded the potential of AI-driven conversational agents. However, creating agents tailored to individual users’ work habits and patterns remains relatively uncharted territory. Although earlier applications of conversational agents in AI productivity tools, as developed by  Grover et al. (2020) and  Kimani et al. (2019), utilized telemetry or sensing data for customization, it has yet to be thoroughly explored, particularly with the latest LLM-based chat agents that offer a more flexible and powerful alternative to their predecessors due to their semantic understanding. Although our work builds upon these existing productivity agent research efforts, several key aspects differentiate our approach. For example, both  Grover et al. (2020) and  Kimani et al. (2019) relied on a custom emotion and context sensing platform, where the collected data was shared with the productivity agent through a secure database. In contrast, our approach leverages modern LLM and readily available workplace telemetry data through Viva Insights [^1] Microsoft’s workplace analytics tool that offers data-driven insights on productivity, well-being, and work habits., eliminating the need to install additional detection systems while maintaining user privacy. This builds on a growing body of work exploring passive sensing and telemetry data for productivity and well-being applications (Nepal et al., 2022, 2023, 2021, 2020, 2024a; Hernandez et al., 2025), although our implementation uniquely combines such data with the semantic understanding capabilities of modern LLMs. Our focus is specifically on facilitating productivity improvement through self-reflection rather than real-time intervention. Recent studies have demonstrated how AI systems can effectively support self-reflection and behavior change across various domains (Bhattacharjee et al., 2024; Kumar et al., 2023; Nepal et al., 2024c). By utilizing GPT-4, our agent can provide more nuanced and contextually relevant responses compared to traditional rule-based conversational agents. We also explicitly compare our agent with alternative presentation modes (dashboard and narrative), providing insights into the relative effectiveness of different approaches. Furthermore, our initial user survey (N=363) provides a broader understanding of user preferences compared to previous studies (e.g., there was no initial user survey done in  Grover et al. (2020), N=70 in  Kimani et al. (2019)). However, these prior systems had some advantages over our approach. For instance,  Grover et al. (2020)’s use of an emotionally expressive avatar might create stronger engagement, while  Kimani et al. (2019)’s focus on well-being through break reminders represents an important aspect that could be incorporated into future versions of our system.
@@
Building on these distinctions, our research aims to conduct a comprehensive two-phase study that examines user preferences and incorporates telemetry data for personalization in AI-driven productivity agents. Our approach goes beyond the existing literature by offering an in-depth analysis of both user preferences and the affordances of telemetry data, which provides valuable insights into user behaviors and work patterns. This combination of large language models (LLMs) with telemetry data enables more accurate and contextually relevant productivity insights than previously possible. It should also be noted that the shift to powerful but proprietary LLMs introduces new challenges in privacy and security  (Das Swain and Saha, 2024; Das Swain et al., 2023a). These include concerns about data retention, model training practices, and the handling of potentially sensitive workplace information. Current approaches to mitigating these risks include using LLMs purely for inference without data retention which might involve working with enterprise versions of LLMs that provide additional privacy guarantees, limiting the granularity of data processed by these models or using locally hosted LLMs (Li et al., 2024; Nepal et al., 2024b). Despite these potential privacy challenges, the implementation of telemetry data with modern LLMs in the context of productivity enhancement remains relatively unexplored. Our study’s innovation lies in the use of advanced LLMs and how we leverage these models to interpret telemetry data, creating a personalized and dynamic productivity tool tailored to individual user needs. We have the position that our findings can pave the way for future research and contribute to the development of more sophisticated productivity tools tailored to the distinct challenges and needs of information workers.